OpenStudy (anonymous):

What is 6th root (723)^3 ?

OpenStudy (anonymous):

\[\sqrt[6]{723}^3\]

jimthompson5910 (jim_thompson5910):

\[\LARGE \sqrt[6]{723^3}\] \[\LARGE 723^{\frac{3}{6}}\] \[\LARGE 723^{\frac{1}{2}}\] \[\LARGE \sqrt{723}\] So \[\LARGE \sqrt[6]{723^3} = \sqrt{723}\]